Is this possible? The probability that a bus will arrive late at the Civic Center is 0.35, and the probability that it will be on time or early is 0.55.

True because 0.35+0.55 <1.

True because 0.35*0.55<1.

False because 0.35+0.55 <1.

False because 0.35 < 0.55.
Answer #1

False because 0.35+0.55 <1.

there will be three condition for the bus

A. either late or

B.early or

C.on time,

all there three event A, B and C are mutually exclusive (or disjoint) event and sum of probability of three event will be 1.

P(A)+(B)+P(C)=1

here P(A)=0.35, P(B)+P( C)=0.55

since P(A)+P(B)+P(C)=0.9 which is not equal to 1. so this is not possible ,

therefore right choice is False because 0.35+0.55 <1.